A 740kg elevator accelerates upward at 1.1m/s2, pulled by a cable of negligible mass. Find the tension force in the cable.
1
Expert's answer
2020-12-17T09:10:17-0500

According to the second Newtons low


"T-mg=ma\to T=m(g+a)=740\cdot(9.8+1.1)=8066(N)"